# NeuroSim V1.5: Improved Software Backbone for Benchmarking Compute-in-Memory Accelerators with Device and Circuit-level Non-idealities

### v1

@@ V Conclusion
    Optimal designs must balance array size, ADC precision, and multi-level cell capability.
@@
    Network complexity influences sensitivity to device-level variations, particularly in larger networks like ResNet-50 and Swin-T.
@@
    Charge-domain computation (e.g., based on nvCap) is a promising device for future large-scale ACIM acceleration.
Looking forward, the emergence of large language models (LLMs) with billions to trillions of parameters presents new challenges beyond on-chip memory capacity. These models will require GB-level high bandwidth memory (HBM) and process-in-memory (PIM) architectures that place compute logic near DRAM dies. While direct LLM and PIM support is beyond NeuroSim V1.5’s current scope, we have initiated efforts toward these capabilities through our work on 3D-stacked DRAM on TPU-like architectures [64, 65].

### v2

@@ V Conclusion and Future Work
    Pareto-optimal designs converge around the 5-7 bit ADC configuration, underscoring the importance of ADC overhead in ACIM architectures.
@@
    Heterogeneous ACIM/DCIM architectures enable transformer acceleration; however, limiting DCIM circuits to ACIM-compatible technology nodes reduces area efficiency. Heterogeneous chiplet integration with separate ACIM and DCIM dies could offer a solution.
@@
    ViTs exhibit lower noise tolerance and area efficiency than CNNs of comparable size. For vision applications where both architectures are viable, CNNs remain better suited for CIM hardware.
Looking forward, several enhancements are planned: lookup table-based noise modeling for complex effects like IR drop, generalizing the PPA framework for flexible system-level exploration similar to CIMLoop, and incorporating time-domain input encoding for broader CIM coverage.
Extending NeuroSim to support large language models is a primary future direction, as transformers are the de-facto architecture for these workloads. LLMs present unique challenges requiring GB-level parameter storage—addressable through multi-die CIM integration or process-in-memory (PIM) architectures with CMOS compute near high-bandwidth memory. We have initiated PIM efforts through our work on 3D-stacked DRAM architectures [69, 70].